OpenStudy (anonymous):

(Can someone please HELP me on this and can you please show the steps.) Reduce all of your answers to simple fractions, mixed numbers, or improper fractions. 1. 2 2/5 x 3 1/3= 2. (3/8) (4/7)= 3. 3 3/5 x 4 4/7= 4. 7/9 x 27/21= 5. 5 x 3/4=

OpenStudy (pfenn1):

Well for 2. (3/8) (4/7) = 12/56=3/14

OpenStudy (pfenn1):

it is pretty straight forward. convert each number to a fraction like 2 2/5=12/5 and perform the multiplication and then reduce the fraction.

OpenStudy (anonymous):

reduce first, multiply last. it is much easier that way

OpenStudy (anonymous):

\[5 \times \frac{ 3}{4}=\frac{5\times 3}{4}=\frac{15}{4}\] for the last one

OpenStudy (anonymous):

if you need mixed number answers \[\frac{15}{4}=3\tfrac{3}{4}\]